OpenStudy (anonymous):

In an effort to reduce her monthly cell phone bill, Hannah talked fifty fewer minutes this month than the previous month. If this represents an eight percent reduction in talk time, how many minutes did Hannah talk on her cell phone the previous month?

OpenStudy (mathstudent55):

A percent problem can be solved as a proportion. 50 minutes are 8% of a larger number. That larger number is the total, or 100%. That is the number we want to know. This is how you state the proportion: 50 minutes is to 8% as x minutes is to 100%.

OpenStudy (anonymous):

In other words 50 * (100/8) right?

OpenStudy (mathstudent55):

Now you write a proportion with fractions: \(\dfrac{50}{8} = \dfrac{x}{100}\) Now you can solve for x using cross multiplication.

OpenStudy (mathstudent55):

@HailKK yes

OpenStudy (anonymous):

Was my answer too straightforward?

OpenStudy (mathstudent55):

Cross multiply: 8x = 50 * 100 8x = 5000 Divide both sides by 8: x = 625

OpenStudy (mathstudent55):

Your answer is correct, but if you are trying to help someone, it is helpful to know how you got to your answer, so the poster can learn from your answer and will be able to solve similar problems without help (as in a test) in the future.

OpenStudy (anonymous):

*face palm* That is called a direct answer. Which is prohibited